OpenStudy (kidthatbro8):

Solve for x: 2/3 (x - 4) = 2x. 2 -2 -8 -4

OpenStudy (anonymous):

x=−103

OpenStudy (kamibug):

Distribute. 2/3x - 8/3 = 2x Subtract 2/3x from both sides. - 8/3 = 4/3x Divide both sides by 4/3. - 2 = x

OpenStudy (kidthatbro8):

alright, thank you @

OpenStudy (kamibug):

No problem! :)